# Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective

### v1

@@ Appendix B Data availability
CIFAR-10 raw data is available from [49] [https://www.cs.toronto.edu/ kriz/cifar.html]. Describable Textures Dataset is available from [50] [https://www.robots.ox.ac.uk/ vgg/data/dtd/]. Mouse brain single-cell ATAC-seq data can be downloaded from Figshare (https://figshare.com/ndownloader/files/25721789). The ATAC-seq datasets have been preprocessed by Luecken et al. ([52]) to characterize gene activities. Mouse embryonic stem cell differentiation data is available in Gene Expression Omnibus with accession code [GSE98664]. The single-cell RNA-seq datasets generated from PBMCs treated with interferon-$\beta$ and from human pancreatic tissues are available from R package Seurat (version 5.0.3) under the name “ifnb” and “panc8”, respectively. Source data are also provided.
Appendix C Code availability
@@
The code for calculating the two diagnosis scores (as an R package), and the code for reproducing the simulation and analysis of this paper are available at https://github.com/zhexuandliu/NE-Reliability-MapContinuity.
Appendix D Acknowledgements
@@
YZ is supported by NSF-DMS grant 2412052 and by the Office of the Vice Chancellor for Research and Graduate Education at the UW Madison with funding from the Wisconsin Alumni Research Foundation. ZL and YZ would like to thank Yixuan Li for suggesting out-of-distribution detection, and thank Sebastien Roch, Zexuan Sun, Xinyu Li and Jingyang Lyu for helpful discussions. RM would like to thank Jonas Fischer, Dmitry Kobak, Stefan Steinerberger and Bin Yu for helpful discussions on t-SNE and UMAP.
@@
Appendix E Supplement to “Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective”
E.1 Derivation of singularity score
@@
Given an input data matrix ${\bm{X}}=[{\bm{x}}_{1},\ldots,{\bm{x}}_{n}]^{\top}$ and its embedding matrix ${\bm{Y}}=[{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}$, without loss of generality, we will derive singularity score for the $n$-th embedding point. Singularity score measures the sensitivity of the embedding point under infinitesimal perturbation. Therefore, we add an infinitesimal perturbation of length $\epsilon$ and direction as a vector $\bm{e}$} of unit length to ${\bm{x}}_{n}$ to analyze the sensitivity of its embedding ${\bm{y}}_{n}$.
@@
We note that ${\bm{y}}_{n}$ is the minimizer of the partial LOO loss involving the $n$-th embedding point:
${\bm{y}}_{n}={\rm argmin}_{{\bm{y}}\in\mathbb{R}^{2}}\sum_{1\leq i\leq n-1}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})$
@@
and denote the perturbed dataset as $\widetilde{{\bm{X}}}$ and $\widetilde{{\bm{y}}}_{n}={\bm{f}}_{n}({\bm{x}}_{n}+\epsilon{\bm{e}})$, which is the mapping of perturbed ${\bm{x}}_{n}$ by LOO-map, we have
$\widetilde{{\bm{y}}}_{n}={\rm argmin}_{{\bm{y}}\in\mathbb{R}^{2}}\sum_{1\leq i\leq n-1}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}(\widetilde{{\bm{X}}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top}).$
@@
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}}\Big)+\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\Bigg]\right|_{{\bm{y}}={\bm{y}}_{n}}=\mathbf{0},$ (S1)
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}(\widetilde{{\bm{X}}}))}{\partial{\bm{y}}}\Big)+\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\Bigg]\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}=\mathbf{0}.$ (S2)
@@
$\left.\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}(\widetilde{{\bm{X}}}))}{\partial{\bm{y}}}\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$
$=$ $\left.\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}}\right|_{{\bm{y}}={\bm{y}}_{n}}+\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}(\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n})$
$+\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v)}{\partial{\bm{y}}\partial v}\right|_{{\bm{y}}={\bm{y}}_{n},v=v_{i,n}({\bm{X}})}(v_{i,n}(\widetilde{{\bm{X}}})-v_{i,n}({\bm{X}}))$
$+o(\|\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n}\|+\|v_{i,n}(\widetilde{{\bm{X}}})-v_{i,n}({\bm{X}})\|).$
Then by doing Taylor’s expansion to $\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$ on $\widetilde{{\bm{y}}}_{n}$, we have
$\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$
$=$ $\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\right|_{{\bm{y}}={\bm{y}}_{n}}+\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}(\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n})+o(\|\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n}\|).$
@@
$\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n}$ $=\epsilon\Big(\frac{\partial{\bm{y}}_{n}}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}+o(\epsilon),$
$v_{i,n}(\widetilde{{\bm{X}}})-v_{i,n}({\bm{X}})$ $=\epsilon\Big(\frac{\partial v_{i,n}({\bm{X}})}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}+o(\epsilon).$
@@
$\left.\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}(\widetilde{{\bm{X}}}))}{\partial{\bm{y}}}\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$ (S3)
$=$ $\left.\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}}\right|_{{\bm{y}}={\bm{y}}_{n}}+\epsilon\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}\Big(\frac{\partial{\bm{y}}_{n}}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}$
$+\epsilon\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v)}{\partial{\bm{y}}\partial v}\right|_{{\bm{y}}={\bm{y}}_{n},v=v_{i,n}({\bm{X}})}\Big(\frac{\partial v_{i,n}({\bm{X}})}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}+o(\epsilon)$
@@
$\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$ (S4)
$=$ $\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\right|_{{\bm{y}}={\bm{y}}_{n}}+\epsilon\left.\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}\Big(\frac{\partial{\bm{y}}_{n}}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}+o(\epsilon).$
@@
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}(\widetilde{{\bm{X}}}))}{\partial{\bm{y}}}\Big)+\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\Bigg]\right|_{{\bm{y}}=\widetilde{{\bm{y}}}_{n}}$
$=$ $\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))}{\partial{\bm{y}}}\Big)+\frac{\partial Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})}{\partial{\bm{y}}}\Bigg]\right|_{{\bm{y}}={\bm{y}}_{n}}$
$+\epsilon\left.\frac{\partial^{2}[\sum_{1\leq i\leq n-1}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})]}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}\Big(\frac{\partial{\bm{y}}_{n}}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}$
$+\epsilon\sum_{1\leq i\leq n-1}\Big[\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v)}{\partial{\bm{y}}\partial v}\right|_{{\bm{y}}={\bm{y}}_{n},v=v_{i,n}({\bm{X}})}\Big(\frac{\partial v_{i,n}({\bm{X}})}{\partial{\bm{x}}_{n}}\Big)^{\top}\Big]{\bm{e}}+o(\epsilon).$
@@
$\frac{\partial^{2}[\sum_{1\leq i\leq n-1}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v_{i,n}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}}]^{\top})]}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}=\frac{\partial^{2}\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}};{\bm{X}})}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}$
@@
$\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n};{\bm{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w(\bm{y}_{i},\bm{y}_{j});v_{i,j}({\bm{X}}))+Z([{\bm{y}}_{1},\ldots,{\bm{y}}_{n}]^{\top}).$
Plug in the first order condition Eqn. S1 and Eqn. S2, we have the change of ${\bm{y}}_{n}$ after the infinitesimal perturbation as
$\widetilde{{\bm{y}}}_{n}-{\bm{y}}_{n}$ $=\epsilon\Big(\frac{\partial{\bm{y}}_{n}}{\partial{\bm{x}}_{n}}\Big)^{\top}{\bm{e}}+o(\epsilon)$
$=-\epsilon{\bm{H}}_{n}^{-1}\sum_{1\leq i\leq n-1}\Big[\left.\frac{\partial^{2}\mathcal{L}(w({\bm{y}}_{i},\bm{y});v)}{\partial{\bm{y}}\partial v}\right|_{{\bm{y}}={\bm{y}}_{n},v=v_{i,n}({\bm{X}})}\Big(\frac{\partial v_{i,n}({\bm{X}})}{\partial{\bm{x}}_{n}}\Big)^{\top}\Big]{\bm{e}}+o(\epsilon)$
where ${\bm{H}}_{n}$ denote the Hessian matrix of the loss function with respect to ${\bm{y}}_{n}$:
${\bm{H}}_{n}=\left.\frac{\partial^{2}\mathfrak{L}({\bm{y}}_{1},\ldots,{\bm{y}}_{n-1},{\bm{y}};{\bm{X}})}{\partial{\bm{y}}\partial{\bm{y}}^{\top}}\right|_{{\bm{y}}={\bm{y}}_{n}}.$
Importantly, ${\bm{H}}_{n}$ is independent of the perturbation direction $\bm{e}$}. A singular Hessian matrix results in the most extreme local discontinuity. The more singular ${\bm{H}}_{n}$ is, the more sensitive the embedding of ${\bm{x}}_{n}$ becomes to infinitesimal perturbations. This is why we define the singularity score for the $n$-th data point as $\lambda_{\min}^{-1}({\bm{H}}_{n})$. The higher the singularity score, the more singular ${\bm{H}}_{n}$ is, and the greater the sensitivity of the embedding of ${\bm{x}}_{n}$ to infinitesimal perturbations.
E.2 Singularity scores for t-SNE, UMAP, LargeVis

### v2

@@ Methods
CIFAR-10 raw data is available from [50] [https://www.cs.toronto.edu/ kriz/cifar.html]. Describable Textures Dataset is available from [51] [https://www.robots.ox.ac.uk/ vgg/data/dtd/]. The pretrained ResNet-18 model is available at [https://huggingface.co/edadaltocg/resnet18_cifar10]. Mouse brain single-cell ATAC-seq data can be downloaded from Figshare [https://figshare.com/ndownloader/files/25721789]. The ATAC-seq datasets have been preprocessed by Luecken et al. ([59]) to characterize gene activities. Mouse embryonic stem cell differentiation data is available in Gene Expression Omnibus with accession code [GSE98664]. The single-cell RNA-seq dataset generated from PBMCs treated with interferon-$\beta$ is available from the R package Seurat (version 5.0.3) under the name ifnb. The single-cell RNA-seq data of human pancreatic tissues is available from the smartseq2 dataset in the R package Seurat (version 5.0.3) under the name panc8. Mouse mammary epithelial single-cell data is available from the R package scRNAseq (version 2.20.0) under the name BachMammaryData. Source data are also provided.
Code availability
@@
The code for calculating the two diagnostic scores (as an R package), and the code for reproducing the simulation and analysis of this paper are available at https://github.com/zhexuandliu/MapContinuity-NE-Reliability.
Acknowledgements
@@
Y.Z. is supported by NSF-DMS grant 2412052 and by the Office of the Vice Chancellor for Research and Graduate Education at the UW Madison with funding from the Wisconsin Alumni Research Foundation. Z.L. and Y.Z. would like to thank Yixuan Li for suggesting out-of-distribution detection, and thank Sebastien Roch, Zexuan Sun, Xinyu Li and Jingyang Lyu for helpful discussions. R.M. would like to thank Jonas Fischer, Dmitry Kobak, Stefan Steinerberger and Bin Yu for helpful discussions on t-SNE and UMAP.
Author contributions
Y.Z. and R.M. conceived the study. Z.L. designed and implemented the method with input from Y.Z. and R.M. Z.L. contributed to the numerical analysis and software implementation. Y.Z. and R.M. designed and developed the theoretical results for the study. Z.L. prepared a draft of the manuscript. Y.Z. and R.M. edited the manuscript.
Competing interests
The Authors declare no competing interests.
@@
Appendix A Supplement to “Assessing and improving reliability of neighbor embedding methods: a map-continuity perspective”
A.1 Derivation of singularity score
@@
Given an input data matrix ${\mathbf{X}}=[{\mathbf{x}}_{1},\ldots,{\mathbf{x}}_{n}]^{\top}$ and its embedding matrix ${\mathbf{Y}}=[{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}$, without loss of generality, we will derive singularity score for the $n$-th embedding point. Singularity score measures the sensitivity of the embedding point under infinitesimal perturbation. Therefore, we add an infinitesimal perturbation of length $\epsilon$ and direction as a vector $\mathbf{e}$} of unit length to ${\mathbf{x}}_{n}$ to analyze the sensitivity of its embedding ${\mathbf{y}}_{n}$.
@@
We note that ${\mathbf{y}}_{n}$ is the minimizer of the partial LOO loss involving the $n$-th embedding point:
${\mathbf{y}}_{n}={\rm argmin}_{{\mathbf{y}}\in\mathbb{R}^{2}}\sum_{1\leq i\leq n-1}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})$
@@
and denote the perturbed dataset as $\widetilde{{\mathbf{X}}}$ and $\widetilde{{\mathbf{y}}}_{n}={\mathbf{f}}_{n}({\mathbf{x}}_{n}+\epsilon{\mathbf{e}})$, which is the mapping of perturbed ${\mathbf{x}}_{n}$ by LOO-map, we have
$\widetilde{{\mathbf{y}}}_{n}={\rm argmin}_{{\mathbf{y}}\in\mathbb{R}^{2}}\sum_{1\leq i\leq n-1}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}(\widetilde{{\mathbf{X}}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top}).$
@@
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}}\Big)+\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\Bigg]\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}=\mathbf{0},$ (S1)
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}(\widetilde{{\mathbf{X}}}))}{\partial{\mathbf{y}}}\Big)+\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\Bigg]\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}=\mathbf{0}.$ (S2)
@@
$\left.\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}(\widetilde{{\mathbf{X}}}))}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$
$=$ $\left.\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}+\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}(\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n})$
$+\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v)}{\partial{\mathbf{y}}\partial v}\right|_{{\mathbf{y}}={\mathbf{y}}_{n},v=v_{i,n}({\mathbf{X}})}(v_{i,n}(\widetilde{{\mathbf{X}}})-v_{i,n}({\mathbf{X}}))$
$+o(\|\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n}\|+\|v_{i,n}(\widetilde{{\mathbf{X}}})-v_{i,n}({\mathbf{X}})\|).$
Then by doing Taylor’s expansion to $\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$ on $\widetilde{{\mathbf{y}}}_{n}$, we have
$\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$
$=$ $\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}+\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}(\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n})+o(\|\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n}\|).$
@@
$\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n}$ $=\epsilon\Big(\frac{\partial{\mathbf{y}}_{n}}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}+o(\epsilon),$
$v_{i,n}(\widetilde{{\mathbf{X}}})-v_{i,n}({\mathbf{X}})$ $=\epsilon\Big(\frac{\partial v_{i,n}({\mathbf{X}})}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}+o(\epsilon).$
@@
$\left.\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}(\widetilde{{\mathbf{X}}}))}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$ (S3)
$=$ $\left.\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}+\epsilon\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}\Big(\frac{\partial{\mathbf{y}}_{n}}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}$
$+\epsilon\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v)}{\partial{\mathbf{y}}\partial v}\right|_{{\mathbf{y}}={\mathbf{y}}_{n},v=v_{i,n}({\mathbf{X}})}\Big(\frac{\partial v_{i,n}({\mathbf{X}})}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}+o(\epsilon)$
@@
$\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$ (S4)
$=$ $\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}+\epsilon\left.\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}\Big(\frac{\partial{\mathbf{y}}_{n}}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}+o(\epsilon).$
@@
$\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}(\widetilde{{\mathbf{X}}}))}{\partial{\mathbf{y}}}\Big)+\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\Bigg]\right|_{{\mathbf{y}}=\widetilde{{\mathbf{y}}}_{n}}$
$=$ $\left.\Bigg[\sum_{1\leq i\leq n-1}\Big(\frac{\partial\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))}{\partial{\mathbf{y}}}\Big)+\frac{\partial Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})}{\partial{\mathbf{y}}}\Bigg]\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}$
$+\epsilon\left.\frac{\partial^{2}[\sum_{1\leq i\leq n-1}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})]}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}\Big(\frac{\partial{\mathbf{y}}_{n}}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}$
$+\epsilon\sum_{1\leq i\leq n-1}\Big[\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v)}{\partial{\mathbf{y}}\partial v}\right|_{{\mathbf{y}}={\mathbf{y}}_{n},v=v_{i,n}({\mathbf{X}})}\Big(\frac{\partial v_{i,n}({\mathbf{X}})}{\partial{\mathbf{x}}_{n}}\Big)^{\top}\Big]{\mathbf{e}}+o(\epsilon).$
@@
$\frac{\partial^{2}[\sum_{1\leq i\leq n-1}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v_{i,n}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}}]^{\top})]}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}=\frac{\partial^{2}\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}};{\mathbf{X}})}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}$
@@
$\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n};{\mathbf{X}})=\sum_{1\leq i<j\leq n}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}}_{j});v_{i,j}({\mathbf{X}}))+Z([{\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n}]^{\top}).$
Plug in the first order condition Eqn. S1 and Eqn. S2, we have the change of ${\mathbf{y}}_{n}$ after the infinitesimal perturbation as
$\widetilde{{\mathbf{y}}}_{n}-{\mathbf{y}}_{n}$ $=\epsilon\Big(\frac{\partial{\mathbf{y}}_{n}}{\partial{\mathbf{x}}_{n}}\Big)^{\top}{\mathbf{e}}+o(\epsilon)$
$=-\epsilon{\mathbf{H}}_{n}^{-1}\sum_{1\leq i\leq n-1}\Big[\left.\frac{\partial^{2}\mathcal{L}(w({\mathbf{y}}_{i},{\mathbf{y}});v)}{\partial{\mathbf{y}}\partial v}\right|_{{\mathbf{y}}={\mathbf{y}}_{n},v=v_{i,n}({\mathbf{X}})}\Big(\frac{\partial v_{i,n}({\mathbf{X}})}{\partial{\mathbf{x}}_{n}}\Big)^{\top}\Big]{\mathbf{e}}+o(\epsilon)$
where ${\mathbf{H}}_{n}$ denote the Hessian matrix of the loss function with respect to ${\mathbf{y}}_{n}$:
${\mathbf{H}}_{n}=\left.\frac{\partial^{2}\mathfrak{L}({\mathbf{y}}_{1},\ldots,{\mathbf{y}}_{n-1},{\mathbf{y}};{\mathbf{X}})}{\partial{\mathbf{y}}\partial{\mathbf{y}}^{\top}}\right|_{{\mathbf{y}}={\mathbf{y}}_{n}}.$
Importantly, ${\mathbf{H}}_{n}$ is independent of the perturbation direction $\mathbf{e}$}. A singular Hessian matrix results in the most extreme local discontinuity. The more singular ${\mathbf{H}}_{n}$ is, the more sensitive the embedding of ${\mathbf{x}}_{n}$ becomes to infinitesimal perturbations. This is why we define the singularity score for the $n$-th data point as $\lambda_{\min}^{-1}({\mathbf{H}}_{n})$. The higher the singularity score, the more singular ${\mathbf{H}}_{n}$ is, and the greater the sensitivity of the embedding of ${\mathbf{x}}_{n}$ to infinitesimal perturbations.
A.2 Singularity scores for t-SNE, UMAP, LargeVis